Calculate the pH of the solution after the addition of each of the given amounts of 0.0502 M HNO3 to a 70.0 mL solution of 0.0750 M aziridine. The p?a of aziridinium is 8.04.

What is the pH of the solution after the addition of 7.78 mL HNO3?

What is the pH of the solution after the addition of 101 mL HNO3?

What is the pH of the solution after the addition of a volume of HNO3 equal to the equivalence point?

What is the pH of the solution after the addition of 108.89 mL HNO3?
Answer #1

Given that concentration of HNO3= 0.0502 M concentration of Aziridine (CHN) = 0.0750 M Volume of C H N = 70.0 mL pka of aziri

b) Addition of 7.78 mL of HNO3: Number of moles of C H N = M*V = 0.0750 M* 70.0 mL = 5.25 mmol Number of moles of HNO3 = M*V

c) Addition of 101 mL of HNO3: Number of moles of C H N = M*V = 0.0750 M* 70.0 mL = 5.25 mmol Number of moles of HNO3 = M*V =

d) At equivalence point Number of moles of C H N = M*V = 0.0750 M* 70.0 mL = 5.25 mmol At the equivalence point, the number o

x.x <a -=10-8.04 0.0301-X x is small, so 0.0301- x = 0.0301 x.x 0.0301 = 9.12011*10-4 x = 2.74x10-10 >x=1.66x10 :: [H*]=1.66x